QUESTION 10


  1. Sulfur dioxide gas reacts with oxygen gas and water according to the chemical reaction below. When 4.5 g of SO2 are mixed with excess O2 and H2O, how many grams of H2SO4 are produced?





1
Expert's answer
2020-09-01T07:37:15-0400

"2SO_2+O_2+2H_2O\to 2H_2SO_4"

Given excess of "O_2" and "H_2O" ,

"2" moles of "SO_2" gives "2" moles of "H_2SO_4" .

Hence,number of moles of "H_2SO_4" produced is equal to number of moles of "SO_2" consumed .

Mass of "SO_2" consumed "=4.5 \ g"

Molar mass of "SO_2=64\ g"

Molar mass of "H_2SO_4=98\ g"

Moles of"\ SO_2" consumed "=\frac{4.5}{64}=" Moles of "H_2SO_4" produced.

Mass of "H_2SO_4" produced "=\frac{4.5}{64}\times 98=6.89\ g"